# A Bilevel Optimization Framework for Imbalanced Data Classification

### v2

@@ 7 Acknowledgments
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education (ORISE) for the DOE under contract number DE-SC0014664. This work was also supported by the Applied Mathematics activity within the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357. We also acknowledge the support by the U.S. Department of Energy for the SciDAC 5 RAPIDS institute
Additionally, this research was supported in part by an appointment with the National Science Foundation (NSF) Mathematical Sciences Graduate Internship (MSGI) Program. This program is administered by ORISE through an interagency agreement between the DOE and NSF. ORISE is managed for the DOE by Oak Ridge Associated Universities (ORAU). All opinions expressed in this paper are the authors’ and do not necessarily reflect the policies and views of NSF, ORAU/ORISE or DOE.

### v3

@@ 7 Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research program under contract number DE-SC0014664. This work was also supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) Program through the FASTMath Institute, RAPIDS institute, distributed resilience systems for science program and the integrated computational and data infrastructure for science discovery program under Contract No. DE-AC02-06CH11357.